3. You perform an experiment to determine the effects of temperature on development in an insect species. You do so by incubating randomly selected eggs and larvae at 4 different and constant temperatures and measure the length (in mm) of the third leg after the last immature ecdysis event, with results shown below (each cell represents an individual insect). Perform a one-way ANOVA on your results (/10) Tema A Temp BTemp. Temo D 12.4 14.5 12.8 13.610.5 0.5 8.7 10.1 9.6 8.1 8 8.4 7.9 7.9 8.7 8.5 8 12.1 10.48.4
Record results of your 1-ANOVA below. Source of Sum of Degrees of Mean ofFvde p-value F-crit variation squares freedom squares Between groups (df) ( (MS) 3 25 412133 4. 323.124 3.2 Within groupsS 0. 3217 Total35 11 . Is there any significant difference between any of the limb lengths in insects reared at different temperatures? In a real experiment, would you have to perform a follow-up Tukey test? .Regardless of whether or not you would have to perform a follow-up Tukey test, perform the test anyway and record the interactions indicated in the table below. Comparison Absolute difference Critical dlifference Significant? A and B A and C B and C Yes /No Yes / No Yes / No 2
Answer #1

One-way ANOVA: C4 versus C5

Source DF SS MS F P
C5 3 77.946 25.982 79.27 0.000
Error 16 5.244 0.328
Total 19 83.190

S = 0.5725 R-Sq = 93.70% R-Sq(adj) = 92.51%


Individual 95% CIs For Mean Based on
Pooled StDev
Level N Mean StDev ---------+---------+---------+---------+
Temp A 5 13.080 0.973 (--*---)
Temp B 5 10.220 0.383 (--*---)
Temp C 5 8.320 0.277 (--*---)
Temp D 5 8.200 0.374 (---*--)

Since p-value=0.000<0.05 hence there is at least one significant difference between the limb lengths in insects reared at different temperature.

So we want to perform Tukey test.

Critical difference-q0.025.4.16 V/ MSE/5= 1.0373 MSE 0.328, qo.o254 4.05 Comparison Absolute dif ference Critical dif ference Significant 025,4.16 /MSE/5-1.0373 A and B 13.080- 10.220 2.86 A and C 13.080 8.320-4.76 A and D 13.080- 8.2004.88 B and C 10.220-8.320 = 1.9 B and D 10.220 8.200 2.02 C and D 8.320 8.2000.12 1.0373 1.0373 1.0373 1.0373 1.0373 1.0373 es Yes es Yes es